Can you please help me solve 5+2(n+1)=2n
Found 2 solutions by stanbon, user_dude2008:
Answer by stanbon(75887)

Can you please help me solve
5+2(n+1)=2n
5 + 2n + 2 = 2n
Subtract 2n from both sides and you will get:
7 = 0
==============
That means there is no solution for the original equation.
=================
Cheers,
Stan H.

Answer by user_dude2008(1862)

5+2(n+1)=2n
5+2n+2=2n
7+2n=2n
7=2n-2n
7=0
Ans: no solution